Are these parallel x+11=y and y-x=-7 What is the slope of the line x+11=y
Answer by checkley77(12844)

The line equation is: Y=mX+b where m=slope
x+11=y
y=x+11 has a slope=1.
y-x=-7
y=x-7 has a slope=1
Therefore these 2 lines are parallel.